# LetsTalk: Latent Diffusion Transformer for Talking Video Synthesis

### v1

@@ 3.3 Multimodal Fusion Schemes for Portrait and Speech Audio
  • Direct Fusion (Fig. 3(a)) directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention.
@@
  • Siamese Fusion (Fig. 3(b)) feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
@@
  • Symbiotic Fusion (Fig. 3(c)) concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
After defining these schemes with fusion compactness ranging from shallow to deep, we aim to adopt reasonable solutions for input modalities (i.e. the reference image of portrait and audio).
@@
Given the reference image, we require that the character of the generated video is consistent with the reference, i.e. using the reference image to strongly guide video generation. Moreover, Since the image modality is close to the video, Symbiotic Fusion is appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ (i.e. frames) times and concatenate it with the latent feature $\bm{F}$ along the spatial dimension [^4] To concatenate latent feature $\bm{F}\in\mathbb{R}^{F\times P\times d}$ with the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$, we extend and repeat the portrait feature along the temporal dimension as $\tilde{\bm{P}}\in\mathbb{R}^{F\times P\times d}$, the shape of concatenated latent feature is $[F,2P,d]$. and feed them into the transformer backbone. In contrast, we observe that Direct Fusion attempts to use shallow portrait feature $\bm{P}$ to guide the learning of all blocks, leading to misaligning the features in the deeper blocks. While Siamese Fusion addresses this misalignment and achieves comparable results to Symbiotic Fusion, it needs to maintain a siamese transformer similar to the backbone, which introduces additional model parameters and complicates the learning task.
Datasets Raw Filtered
#IDs #Hours #IDs #Hours
HDTF 411 15.75 356 13.95
CelebV-HQ 35,666 68 22,423 41.27
Train - - 20,501 49.69
Eval - - 2,278 5.522
Figure 4: The dataset splits for HDTF and CelebV-HQ, including raw/filtered data and training/evaluation distributions.
@@
As for the audio, we require it to reasonably drive the animation, especially the lip motion [52]. However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Therefore, we aim for the audio to have a minimal correlation with these animations. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow [47] and extract richer semantic information from different layers of the wav2vec encoder [36]. A sliding window mechanism is employed to capture long-term sequential audio information. Through nonlinear mapping, we obtain audio tokens enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
4 Experiment
4.1 Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF [53] and CelebV-HQ [56]. These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in [47]. We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 22,423 clips with 41.27 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs [23]. SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method [12] and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) [30], SadTalker (138M) [52], AniPortrait (2.8B) [45], and Hallo (2.4B) [47]. The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes, as shown in Fig. 4.
@@ 4.1 Experimental Setups
Figure 5: The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
@@
Figure 6: The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
@@
Ours (Large) 37.134 219.119 0.632 1.962
Table 1: The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
@@
Table 2: The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
Implementation details Our approach differs from previous methods [41, 47, 45] that require a two-stage training process. We adopt an end-to-end training strategy. Thanks to the scalability of transformers, we set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability. Following common practices in generative modeling works [31, 2], an exponential moving average (EMA) weight with a decay rate of $0.9999$ is maintained throughout training, and all reported results are directly derived from the EMA model. Additionally, we utilize a pre-trained VAE from Stable Diffusion 1.4, which remains frozen during training. Please refer to Sec. 7 for more implementation details.
4.2 Quantitative Results
4.3 Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 9) and the CelebV-HQ dataset (Fig. 10). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment. More results can be found in Sec. 8.
4.4 Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 3 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training. We conduct experiments on the HDTF dataset using the LetsTalk-B model to validate the impact of different fusion schemes.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues [52] (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 4 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation. To validate the impact of different fusion schemes on diversity, we serve LetsTalk as an unconditional video generator [^5] Filling audio embeddings with 00 for fusion. and compare the diversity of character actions in the generated videos. The results in Fig. 7 demonstrate that the deeper the fusion compactness, the more static the generated video. The tightest Symbiotic Fusion generates a portrait with almost no motions, Siamese Fusion produces slight motions, while the least tight Direct Fusion we adopt results in natural motions.
Figure 7: Portrait animation results without audio guidance for three fusion schemes. Siamese and Symbiotic fusion schemes with high fusion compactness limit the diversity of video generation, while Direct fusion preserves this diversity.
@@ 4.4 Ablation Study
Table 3: Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
@@
Table 4: Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
5 Conclusion
We present LetsTalk, a novel multimodality-guided diffusion transformer to achieve superior audio-driven portrait animation results using audio input. In LetsTalk, we introduce modular spatial-temporal attention mechanisms into our transformer backbone to enhance temporal consistency. We summarize three fusion schemes to handle multimodality and deeply analyze their impact. We then adopt the optimal fusion schemes for input reference portrait and audio, respectively. Our extensive experiments demonstrate the effectiveness of LetsTalk in achieving state-of-the-art portrait animation results with a good balance between performance and payload.
Supplementary Material
6 Network Architecture
6.1 Backbone
In LetsTalk, we borrow the Variational AutoEncoder (VAE) [25] from Stable Diffusion [34]. Given the reference portrait image and noisy video, we first encode them using the encoder in VAE. Then we patchify and flatten them along the spatial dimension. Finally, we use an MLP to project them into the latent space. The embeddings in the latent space will be fed into our transformer backbone. Specifically, our LetsTalk model is constructed using a series of N transformer blocks, each with a hidden dimension of D and H multi-head attention heads. Inspired by ViT and DiT, we explore two configurations of LetsTalk with varying numbers of parameters, while maintaining a fixed patch size of 2, as detailed in  Tab. 5. Via the encoder in VAE, the downsampling factor is $8$, the input reference portrait image $\bm{I}_{\text{p}}\in\mathbb{R}^{256\times 256\times 3}$ and the noisy video $\bm{V}\in\mathbb{R}^{16\times 256\times 256\times 3}$ are encoded as a portrait feature $\bm{P}\in\mathbb{R}^{32\times 32\times 4}$ and a video feature $\bm{F}\in\mathbb{R}^{16\times 32\times 32\times 4}$. Next, the two features are partitioned according to $2\times 2$ patch size and flattened as $\bm{P}\in\mathbb{R}^{256\times 16}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 16}$. Take our base model (i.e. LetsTalk-Base) as an example, we use an MLP to project both features into the latent space as $\bm{P}\in\mathbb{R}^{256\times 768}$ and $\bm{F}\in\mathbb{R}^{16\times 256\times 768}$.
In the spatial-temporal-aware block, we integrate the block details into our pipeline to model the relationship between temporal and spatial attention. Inspired by  [28], we introduce a fully connected layer between the temporal and spatial attention mechanisms. This layer applies a linear transformation to the output of the temporal attention, enhancing the learning of temporal features and ensuring consistent dimensionality for subsequent processing.

### v2

@@ 3.2. Analysis of Multimodal Fusion Schemes
  • Direct Fusion directly feeds the condition modality into each block of the backbone and fuses it with the inner features via cross-attention;
@@
  • Siamese Fusion feeds the condition modality into another similar transformer, extracts features from each block, and fuses them with the features in the corresponding block of the backbone;
@@
  • Symbiotic Fusion concatenates condition modality with inputs before feeding it into the backbone and achieving seamless fusion via inherent self-attention.
Integration methods: The choice between self-attention and cross-attention involves trade-offs in computational efficiency and optimization. Cross-attention, with smaller interaction matrices, is computationally lighter and easier to optimize, especially in limited data scenarios, as demonstrated in early text-to-image models (Rombach et al., 2022; Nichol et al., 2021). In contrast, self-attention, which requires larger matrices for multimodal token alignment, is more challenging to optimize but achieves higher embedding quality with sufficient data (Wang et al., 2024c). Alternatively, addition, which sums hidden states, provides a simpler and more computationally efficient approach, particularly useful when conditional inputs aim to highlight specific local features in the latent space.
Condition States: The static versus dynamic nature of conditions reflects their constraint strength. Dynamic states, as in Siamese and Symbiotic Fusion, enable hierarchical refinement of conditional features during training, aligning with findings from ReferenceNet (Hu, 2024) and ControlNet (Zhang et al., 2023b) that multi-level conditioning enhances generation controllability. This flexibility improves feature alignment and model performance.
Parameters Sharing: Siamese Fusion uses separate network parameters for conditions and latent noise, while Symbiotic Fusion shares the same parameters. Direct Attention, relying solely on cross-attention, also does not share parameters.
Modality Adaptation: The adaptability of fusion schemes to different modalities is key to their success in multimodal tasks. Direct Fusion, though offering weak alignment, is highly cost-effective for modalities with significant differences, such as text and image, making it ideal for resource-constrained scenarios. Siamese Fusion offers enhanced design flexibility by allowing adjustable modality alignment. This is exemplified in models such as ControlNet (Zhang et al., 2023b), which employs visual cue-level alignment, and AnimateAnyone (Hu, 2024), which utilizes pixel-level alignment. The strength of alignment can thus be precisely tailored to meet the specific requirements of the task at hand. Symbiotic Fusion enhances parameter efficiency compared to Siamese Fusion, achieving strong alignment with fewer parameters. This makes it particularly effective for large-scale applications, such as AnimateAnyone (Hu, 2024), where shared parameters ensure robust and efficient multimodal integration. The comparative analysis of the three fusion approaches are summarized in Tab. 1
3.3. Guidance on Reference Portrait and Audio
@@
Given the reference image, we require that the character of the generated video is consistent with the reference. Since the image modality is close to the video, Symbiotic Fusion proves appropriate. Specifically, we first encode the reference image and obtain the portrait feature $\bm{P}\in\mathbb{R}^{P\times d}$ in the latent space. Then we temporally repeat the portrait feature $\bm{P}$ with $F$ frames and concatenate it with the latent feature $\bm{Z_{s}}\in\mathbb{R}^{F\times P\times d}$ along the spatial dimension, which can be formulated as: $\bm{Z}=\text{concat}\left(\,\bm{Z_{s}},\,\tilde{\bm{P}}\,\right)$ denotes the temporally extended portrait feature. This operation yields a concatenated latent feature $\bm{Z}\in\mathbb{R}^{F\times 2P\times d}$. By leveraging shared self-attention in the fused feature space, this method maintains consistent character representation across frames. In contrast, Direct Fusion uses shallow portrait features, leading to misalignment in deeper layers. While Siamese Fusion mitigates this by maintaining parallel transformer branches, it increases computational complexity without significant advantages over Symbiotic Fusion.
@@
As for the audio, we require it to reasonably drive the animation, especially the lip motion (Zhang et al., 2023a). However, certain unconscious actions (e.g. blinking and head movements) are almost independent of the audio yet serve to diversity. Further considering that the modality of audio is quite different from the video, using Siamese Fusion and Symbiotic Fusion often causes ambiguity in the backbone learning, damaging the diversity. In a nutshell, Direct Fusion is concise yet effective for fusing the input audio. In addition, for the audio representation, to accommodate the tokenization mechanism of the transformer, we follow (Xu et al., 2024b) and extract richer semantic information from different layers of the wav2vec encoder (Schneider et al., 2019). Through nonlinear mapping, we obtain audio tokens $\bm{C}\in\mathbb{R}^{F\times T\times d^{\prime}}$ enriched with semantic content, which are aligned with latent features in the same feature space. Temporal-level audio representations are then synchronized with video sequences, resulting in the semantic audio representation. The experiments in Sec. 4.4 verify the efficacy of our mechanism in feature fusion.
Figure 4. The illustration of the long-duration generation.
3.4. Long-Duration Generation
Unlike the previous frame-by-frame generation approach (Shen et al., 2023; Stypułkowski et al., 2024), our end-to-end diffusion transformer generates a video clip in a single sampling process. In diffusion-based video generation tasks, ensuring proper temporal alignment is critical for maintaining coherence and consistency across generated video sequences. Additionally, due to the enhanced realism and dynamism of the talking videos we generate, direct transitions between consecutive clips are often not feasible. To address this, we introduce a memory bank that stores information from previously generated clips. This data undergoes patchification and temporal transpose operations, transforming it into $\bm{M}\in\mathbb{R}^{P\times f\times d}$, where $f$ represents the number of frames selected from the prior clips. Leveraging the scalability of the Transformer architecture, the memory bank directly provides prior information through Symbiotic Fusion with $\bm{Z_{t}}$. However, the inherent distribution discrepancy between training (real data) and inference (synthetic data) poses a fundamental challenge for diffusion models, with error accumulation becoming increasingly severe as video length grows.
To mitigate this latent risk, we develop a noise-regularized training strategy. Specifically, we first establish that diffusion models inherently retain residual noise even after finite-step denoising due to the discrete approximation of continuous noise schedules, which progressively accumulates as the video sequence extends. Building on the insight, we introduce a bounded stochastic noiser that perturbs the memory bank contents with controlled Gaussian noise $\epsilon\sim\mathcal{N}(0,{\sigma_{mini}}^{2}\mathbf{I})$ (see Fig. 4) into the memory bank during training. Inspired by Hallo2 (Cui et al., 2024), we further apply a patch-drop data augmentation strategy to the memory bank. Unlike Hallo2 which protects face patches from dropping, we employ completely random dropping to enhance the model’s robustness to the memory bank. This deliberate corruption serves two purposes: (1) it faithfully simulates inference-time distribution shifts, and (2) it preserves crucial temporal dependencies. The resulting framework demonstrates significantly improved robustness against error propagation in long-duration generation tasks.
@@ 3.4. Long-Duration Generation
Figure 5. The qualitative comparisons with other cutting-edge methods on the HDTF dataset. Our method achieves better audio-animation alignment (e.g. lip motions) and produces expressive results.
4. Experiment
4.1. Experimental Setups
Datasets We primarily conduct comprehensive experiments on two public datasets: HDTF (Zhang et al., 2021) and CelebV-HQ (Zhu et al., 2022). These datasets include videos that consist of individuals of various ages, ethnicities, and genders, captured in half-body or close-up shots against varied indoor and outdoor backgrounds. To ensure high-quality training data, we adopt the preprocess scheme used in (Xu et al., 2024b). We first exclude videos with scene changes, significant camera movements, excessive facial motion, and side-facing shots. Then we refine the remaining videos and retain only single-person speaking videos with strong audio-lip correspondence. After refinement, we obtain 356 clips with 13.95 hours from the HDTF, and 24438 clips with 44.91 hours from the CelebV-HQ dataset. During training, we extract 16-frame video clips from these datasets using a specific sampling interval, each frame is resized to $256\times 256$ resolution.
Evaluation Metrics The evaluation metrics utilized in the portrait image animation approach include Fréchet Inception Distance (FID), Fréchet Video Distance (FVD), Structural Similarity Index Measure (SSIM) and Expression-FID (E-FID). Specifically, FID and FVD measure the similarity between generated images and real data, with lower values indicating better performance and thus more realistic outputs (Karras et al., 2019). SSIM indices measure the structural similarity between ground truth videos and generated videos. E-FID leverages Inception features to assess image authenticity, providing a more precise fidelity measure. It extracts facial expression parameters via a reconstruction method (Deng et al., 2019) and computes FID to quantify the gap between generated and ground truth expressions.
Baseline In experiments, we compare our method with the publicly available implementations of DreamTalk (34M) (Ma et al., 2023), SadTalker (138M) (Zhang et al., 2023a), AniPortrait (2.8B) (Wei et al., 2024), Hallo (2.4B) (Xu et al., 2024b), Hallo (2.4B) (Cui et al., 2024) and EchoMimic (2.6B) (Chen et al., 2024). The evaluation is performed on the HDTF and CelebV-HQ Dataset, utilizing a training and testing split where $90\%$ of the identity data was allocated for training purposes.
@@ 4.1. Experimental Setups
Figure 6. The qualitative comparisons with the existing portrait image animation approaches on the CelebV-HQ dataset. Our method achieves better portrait consistency.
@@
Ours (Large) 31.761 219.119 0.632 1.962
Table 2. The quantitative comparisons with the existed portrait image animation approaches on the HDTF dataset.
Implementation details Our approach, in contrast to previous methods (Tian et al., 2024; Xu et al., 2024b; Wei et al., 2024) that rely on a two-stage training process, adopts an end-to-end training strategy. We set up models with two sizes: the Base model, which is suitable for simple fine-tuning, and the Large model, which demonstrates better generalization ability.
4.2. Quantitative Results
4.3. Qualitative Results
We have visualized the results of different methods on the HDTF dataset (Fig. 8) and the CelebV-HQ dataset (Fig. 9). From the visualization results, it is evident that our method performs exceptionally well on both datasets. In comparison to the visual cues most strongly associated with audio (such as facial muscle and lip motions), our results closely align with the standard Ground Truth. Benefitting from our Symbiotic Fusion, our generated videos maintain strict portrait consistency. Furthermore, by employing Direct Fusion of audio, the characters exhibit strong expressiveness and diversity in our generated videos while achieving audio-animation alignment.
@@ 4.3. Qualitative Results
Table 3. The quantitative comparisons with the existed portrait image animation approaches on the CelebV-HQ dataset.
@@
Table 4. Experimental results of three multimodal fusion schemes, guided by reference portraits, on the HDTF validation dataset, employing the Letstalk-B model.
@@
Table 5. Experimental results of three multimodal fusion schemes, driven by audio inputs, on the HDTF validation dataset, employing the Letstalk-B model.
4.4. Ablation Study
In this section, we analyze the impact of different fusion schemes as illustrated in Fig. 3 on guidance using reference portrait and audio. Specifically, for analyzing Siamese Fusion, we have to design Siamese transformers for both reference portrait and audio to produce guidance for each block of the backbone. For the reference portrait, we leverage a DiT model pre-trained on ImageNet. For the audio, we define a transformer whose blocks are equipped with two-layer MLPs. The parameters of these transformers are simultaneously updated during training.
Analysis on the Driven Audio Fusion The way we fuse audio affects both audio-animation alignment and diversity. Audio-animation alignment refers to specific visual cues (Zhang et al., 2023a) (e.g. facial muscle and lip motions), whereas the diversity is related to the speech-agnostic motions (e.g. blinking and neck moving) that often appear as unconscious actions. We observe that diversity mainly comes from the transformer backbone during denoising. Thus, we aim to fuse the audio to achieve alignment while preserving the backbone’s inherent diversity. Quantitative results in Tab. 5 indicate that Direct Fusion outperforms the other schemes in establishing the correspondence between audio and animation.
Method FID $\downarrow$ FVD $\downarrow$
Baseline model with Memory Bank 52.243 276.16
+ Regularized Training (w/o patch drop) 32.746 264.709
+ Regularized Training (w patch drop) 34.821 251.684
Table 6. Experimental results of long video generation. Experimental results comparing three approaches for 64-second video generation (100 clips): (1) memory bank baseline, (2) regularized training, and (3) patch drop augmentation.
4.5. Long-Duration Generation
To validate the effectiveness of our proposed approach for long video generation, we conducted experiments to analyze the quality of 100 consecutive video clips (about 64 second) generated under three settings: (1) the baseline model (LetsTalk-L) with with the memory bank, (2) the method augmented with regularized training strategy, and (3) the method further enhanced with the patch drop augmentation. Our results in Tab. 6 demonstrate that the memory bank significantly improves temporal consistency, while the regularized training strategy further mitigates degradation in later clips, leading to higher overall video quality. The results demonstrate that regularized training strategy primarily enhances single-frame quality, significantly reducing FID from 52.243 to 32.746, while maintaining temporal consistency (FVD 264.709). On the other hand, patch drop augmentation further improves long-term motion coherence, achieving the best FVD (251.684) with only a slight FID increase (34.821), confirming its complementary role in stabilizing long-duration generation.
5. Conclusion
We introduced LetsTalk, an efficient linear diffusion transformer for high-fidelity talking head synthesis, tackling core challenges in multimodal fusion and long-term temporal coherence. Our approach integrates a deep compression autoencoder for compact latent representation, alongside a systematic exploration of fusion strategies—yielding an optimal framework with Symbiotic Fusion for portrait consistency and Direct Fusion for audio-video synchronization. The memory bank mechanism and noise-regularized training further mitigate error accumulation in extended sequences. Extensive experiments demonstrate LetsTalk’s state-of-the-art performance, generating lifelike, temporally coherent animations with enhanced diversity, all while maintaining superior efficiency through its optimized architecture.
Appendix A Preliminaries
Diffusion Formulation The diffusion process incrementally corrupts data $x_{0}$ with Gaussian noise: $q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}$ are predetermined hyperparameters. By reparameterizing, we obtain: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$, where $\epsilon\sim\mathcal{N}(0,\mathbf{I})$, forming a Markov chain over $T$ steps.
(2) $\begin{split}\mathcal{L}(\theta)=&\sum_{t}\mathcal{D}_{\text{KL}}(q^{*}(x_{t-1}|x_{t},x_{0})\,\|\,p_{\theta}(x_{t-1}|x_{t}))\\ &-p(x_{0}|x_{1}).\end{split}$
(3) $\mathcal{L}_{\text{simple}}(\theta)=\|\epsilon_{\theta}(x)-\epsilon(x)\|_{2}^{2}.$
Diffusion Transformer To better handle the multi-modality fusion in the talking head task, we are motivated to extend the Diffusion Transformer (DiT) in our framework. DiT builds upon the Vision Transformer (ViT) architecture, which processes sequences of image patches. This sequential patch-based processing mechanism natively supports the fusion of multi-modality sufficiently and efficiently. Different from ViT, DiT incorporates a variational autoencoder (VAE) (Kingma, 2013) in the framework, thus runs in the latent space, rather than pixel space. Such design enables DiT to handle the expensive high-resolution generation task, which brings more obvious benefits in the video generation task.